OpenStudy (ohmath123):

Math Puzzle Please HELP! A group of friends decide to rent a house in June Lake for a ski trip. They each had to chip in $70 for the rental. If they had been able to convince 3 more people to go, the cost per person would have been reduced by $14. What was the rent for the week?

rebeccaxhawaii (rebeccaxhawaii):

we don't know how many people they started with x+3

jhonyy9 (jhonyy9):

do you understand what rebecca have wrote above ?

OpenStudy (ohmath123):

yes i did but I think we're suppose to guess that? but how?

jhonyy9 (jhonyy9):

for the first case you can writing 70x = S - let be S the rent cost sum of money /week and after 3 more people you can writing 14(x+3) = S - so now you have got a system af equation with two unknowed terms x and S ,what you need to solve it using substitution method or elimination - what you like - ok. ? 70x=S 14(x+3)=S

OpenStudy (ohmath123):

makes sense, thanks @jhonyy9

jhonyy9 (jhonyy9):

np was my pleasure bye bye